Gomez Corporation purchased land, land improvements, and a building for $480,000. The assets are valued as follows:

Asset Appraised Value
Land $100,000
Land Improvements 75,000
Building 325,000

How much should be debited to each account?

Rolland Company paid $250,000 to acquire a 50 ton press, $11,200 in freight charges, $3,500 for installation and $600 for testing. During installation, a Rolland employee accidentally broke a part that was repaired for $300.

What is the capitalized cost of the press?

$264,700
$265,300
$265,600
$261,200
Answer #1
1
Amount debited to:
Land 96000 =480000/(100000+75000+325000)*100000
Land Improvements 72000 =480000/(100000+75000+325000)*75000
Building 312000 =480000/(100000+75000+325000)*325000
2
Amount paid 250000
Add:
Freight charges 11200
Installation 3500
Testing 600
Capitalized cost of the press 265300